Suppose I am a grandmother. My granddaughter is about as old in days as my daughter is in weeks. She is also as many months old as I am in years. Together, the three of us are 140 years old. How old am I?
Answer by ankor@dixie-net.com(22740)

Suppose I am a grandmother.
let a = grandma's age
let d = daughter's age
let g = granddaughter's age
My granddaughter is about as old in days as my daughter is in weeks.
365g = 52d
divide both sides by 52 Approx
7g = d
She is also as many months old as I am in years.
12g = a
Together, the three of us are 140 years old.
a + d + g = 140
replace d with 7g, replace a with 12g
12g + 7g + g = 140
20g = 140
g = 140/20
g = 7 yrs old is the granddaughter
:
How old am I?
a = 12(7)
a = 84 yrs is grandma